OpenStudy (anonymous):

determine whether the ratio in each pair can form a proporition by writing each ratio in simplest form 15/45 . 3/15

OpenStudy (turingtest):

15/45=1/3 3/15=1/5 i.e. they are not proportional

OpenStudy (anonymous):

how do you know these things? everyone here is sooo smart

OpenStudy (turingtest):

divide the top and bottom by the largest number that goes into both. for 15/45, 15 and 45 can be divided by 15. if you can't see that easily at least divide them both by 5, since all numbers ending in 5 are divisible by 5 15/5 ----- = 3/9 45/4 now you can divide top and bottom by 3 to get 1/3 3/15 is only divisible by 3. Dividing top and bottom by 3 gives 1/5 1/5 does not equal 1/3, therefor they are not proportional.